# ProtoECGNet: Case-Based Interpretable Deep Learning for Multi-Label ECG Classification with Contrastive Learning

### v2

@@ 3.5 Training Procedure
This stage used binary cross-entropy loss for multi-label prediction, with L1 regularization to encourage sparsity in the learned fusion weights. The sparsity constraint promoted more selective use of prototype information. Branch-specific classifiers were trained in the same way, to assess if the fusion classifier yielded performance benefits over simple macro-averaging. Additional training and tuning details are provided in Appendices A and B.
3.6 Manual Evaluation of Prototype Quality
To assess the interpretability and perceived clinical utility of the learned prototypes, we conducted a structured review with two practicing physicians. The evaluation focused on the final ProtoECGNet model trained with the full prototype loss suite, including contrastive regularization. Details on the graphical user interface (GUI) used for clinician review and a screenshot of the interface are provided in Appendix E.
Each criterion was rated on a 1–5 Likert scale. Reviewers were instructed to score each prototype independently based on the projected ECG segment (see Appendices E and F for additional details). No test cases or model predictions were presented. ECGs that were identified as having label errors by the clinicians were excluded. We report the mean scores and 95% confidence intervals across reviewers for each evaluation criterion.
4 Results & Discussion
We evaluate ProtoECGNet in terms of diagnostic classification performance and prototype interpretability. All results are reported on the held-out PTB-XL test set (fold 10). Macro-AUROC is used as the primary evaluation metric throughout to allow comparison to the PTB-XL benchmarking study Strodthoff et al. (2021).
Experiment: To evaluate whether aligning prototype types with clinical reasoning modalities improves performance, we compare models trained on all 71 PTB-XL labels using (1) a single prototype type (e.g., only 1D or 2D CNNs), and (2) the full ProtoECGNet multi-branch model. Each architecture is evaluated as a black-box baseline and as a prototype model with and without contrastive loss. Multi-branch prototype outputs are fused either via simple macro-averaging of branch-specific classifier predictions or a learned fusion classifier on the similarity scores.
Results: The best-performing single-branch model was the 2D partial prototype model with contrastive loss, achieving a macro-AUROC of 0.9137 (See Table 1). However, the full ProtoECGNet model—combining 1D, 2D partial, and 2D global branches—achieved the highest overall performance when using a learned fusion classifier, with a macro-AUROC of 0.9248. This exceeds all single-branch models and closely matches the best-reported performance in the PTB-XL benchmarking study (macro-AUROC 0.925 single-model; 0.929 ensemble) Strodthoff et al. (2021).
Table 1: Effect of contrastive prototype loss on macro-AUROC across branch-specific, single-branch, and multi-branch settings for ProtoECGNet.
Setting Model (Label Set) No Contrastive w/ Contrastive
Branch-Specific Labels Rhythm Branch (16 labels) 0.8730 0.8758
Morphology Branch (52 labels) 0.9023 0.9056
Global Branch (3 labels) 0.7461 0.8606
Full 71-Label, Single Branch 1D Prototype Model 0.8462 0.8794
2D Partial Prototype Model 0.9031 0.9137
2D Global Prototype Model 0.8719 0.8983
Full 71-Label, Multi-Branch Macro Aggregation 0.8918 0.9048
Fusion Classifier 0.9010 0.9248
Discussion: These results support the hypothesis that modeling distinct diagnostic reasoning types using specialized prototype branches can improve performance. While several single-branch models perform strongly—such as the 2D partial prototype model (0.9137)—they apply a uniform prototype structure to all diagnoses, regardless of their underlying interpretive requirements. In contrast, ProtoECGNet’s multi-branch design uses rhythm prototypes for temporal abnormalities, time-localized 2D prototypes for focal morphological findings, and global 2D prototypes for diffuse ECG patterns. The resulting fusion classifier achieves the highest overall macro-AUROC (0.9248), indicating that this domain-tailored architectural structure does not compromise diagnostic performance. Although the performance gains are modest overall, ProtoECGNet matches or exceeds all single-branch models and enables diagnosis-appropriate prototype types that are intended to improve downstream interpretability.
Results: Across all settings, contrastive loss improved performance over the standard prototype loss terms (see Table 1). The largest gains were seen in the 2D global branch, where contrastive loss raised macro-AUROC from 0.7461 to 0.8606 in the branch-specific model and from 0.8719 to 0.8983 in the full-label model. Improvements were also observed in the 2D partial model (0.9031 to 0.9137), 1D model (0.8462 to 0.8794), and both multi-branch models. The contrastive loss fusion classifier achieved the highest score (0.9248)
Discussion: These results confirm that contrastive prototype loss consistently enhances diagnostic performance. Improvements were observed across all architectures and label settings, suggesting that our loss formulation provides generalizable gains. This highlights the importance of class co-occurrence-informed regularization in multi-label prototype learning.
4.3 How does ProtoECGNet compare to black-box baselines?
Experiment: For each model variant—1D, 2D partial, 2D global, and multi-branch—we compare a black-box ResNet baseline with the corresponding prototype models (with and without contrastive loss). This allows us to assess whether our interpretable architecture sacrifices diagnostic performance.
Results: Black-box models achieved strong results—0.9219 for the 1D ResNet18 baseline and 0.9060 for 2D ResNet18 (2D Global in Table 2). Prototype-based models trained with contrastive loss approximately matched or surpassed these: the 2D partial model achieved 0.9137 and the 2D global model reached 0.8983. The full ProtoECGNet fusion classifier outperformed all individual black-box models, achieving 0.9248.
Table 2: Macro-AUROC comparison of black-box baselines vs. prototype-based models trained on all 71 labels.
Model Type Black-Box Proto (No Contrastive) Proto (w/ Contrastive)
1D 0.9219 0.8462 0.8794
2D Partial 0.9031 0.9137
2D Global 0.9060 0.8719 0.8983
Multi-Branch (Macro Agg.) 0.9052 0.8918 0.9048
Multi-Branch (Fusion Classifier) 0.9010 0.9248
Discussion: These results demonstrate that interpretable prototype models can achieve diagnostic performance comparable to traditional black-box CNNs. The 2D partial prototype model surpassed the 2D black-box model, and the final multi-branch model outperformed even the best-performing single-architecture black-box (1D ResNet18). This suggests that prototype-based reasoning, when designed with task-specific structure and contrastive regularization, can achieve state-of-the-art classification while supporting transparent, case-based explanations.
4.4 Are the learned prototypes clinically meaningful?
Experiment: To assess interpretability, we conducted a structured review of all projected prototypes from the final contrastive model. Two physicians—one with board certification in cardiology and one in internal medicine—independently rated each prototype on a 1–5 Likert scale for two criteria: representativeness (how well the prototype exemplifies the assigned class) and clarity (how visually interpretable and artifact-free the segment is).
Reviewer Representativeness (95% CI) Clarity (95% CI)
Cardiologist 4.29 [4.22, 4.35] 4.48 [4.42, 4.54]
Internist 3.59 [3.52, 3.66] 4.73 [4.69, 4.77]
Discussion: These results suggest ProtoECGNet learns meaningful and recognizable ECG patterns which align with clinician expectations. High average scores for both representativeness and clarity indicate that the prototypes are visually interpretable and diagnostically appropriate. This review provides initial validation that the model generates human-interpretable, class-representative prototypes across diverse diagnostic categories but it does not assess downstream utility. Future work plans to include this assessment as a larger user study.
4.5 How does ProtoECGNet support case-based explanations?
4.6 Limitations
This study has several limitations that inform directions for future work. First, our prototype taxonomy relies on a manually defined grouping of PTB-XL diagnostic labels into rhythm, morphology, and global categories. While this structure was clinically motivated and grounded in ECG interpretation heuristics, it introduces simplifications that may not fully capture the nuances of certain diagnoses. Some conditions may exhibit features that span multiple reasoning modalities (e.g., both rhythm and morphology), which are not explicitly modeled in the current architecture. Future work could explore more flexible or data-driven grouping strategies.
Third, we used ResNet-based CNNs as the backbone architecture across all experiments to ensure consistency in comparison. While this design choice was appropriate for isolating the effect of prototype modeling and contrastive loss, it may limit performance relative to more recent architectures. ProtoECGNet is modular by design and compatible with alternative backbones, which should be explored to assess potential gains in both accuracy and interpretability.
Finally, our contrastive prototype loss relies on empirical label co-occurrence statistics derived from the PTB-XL training set. These statistics may reflect dataset-specific artifacts or biases and may not generalize across institutions or populations. Methods that learn co-occurrence-aware regularization in a more adaptive or transferable manner remain an area for future work.
5 Conclusions
Our work demonstrated that prototype-based learning can be scaled to complex, multi-label medical time-series tasks like ECG classification without sacrificing performance. By separating diagnostic labels into rhythm, morphology, and global categories, ProtoECGNet enabled the model to learn representations aligned with clinically distinct reasoning processes. Each branch learned prototypes tailored to its diagnostic task, and predictions were grounded in real ECG segments rather than abstract feature maps. Our proposed contrastive loss improved performance by structuring the prototype space to reflect label co-occurrence patterns observed in real-world data. Across all architectures, contrastive training consistently outperformed standard prototype objectives and black-box baselines. Clinician ratings provided an initial indication that the resulting prototypes were clear and representative of their assigned classes. While our evaluation focused on ECGs, the architectural design and contrastive prototype formulation are broadly applicable to other structured time-series tasks in medicine. Future work will explore prospective clinical validation, integration with electronic health records, and expansion to other modalities.
[…]
This work was funded in part by the National Institutes of Health, specifically grant number R00NS114850 to BKB.
Appendix A Implementation Details
@@
Hyperparameter tuning was conducted using Optuna with 200 trials per model variant. Tuned parameters included learning rate, weight decay, dropout rate, and learning rate scheduler configuration. Both the joint training phase and the classifier stage (including the fusion classifier and branch-specific classifiers) were tuned independently using this strategy. Loss weightings for each prototype loss term ($\lambda_{\text{clst}},\lambda_{\text{sep}},\lambda_{\text{div}},\lambda_{\text{cntrst}}$) were also included as tunable parameters.
When contrastive loss was not used, we adopted the prototype loss coefficients from Barnett et al. (2024). For contrastive models, we conducted a dedicated 1000-trial Optuna sweep to tune all four loss coefficients. This sweep fixed non-loss hyperparameters (dropout = 0.3, batch size = 32, learning rate = 0.001, scheduler = ReduceLROnPlateau, L2 weight decay = $1\times 10^{-4}$), and only optimized the four loss weights. This was run separately for the 1D rhythm, 2D partial, and 2D global prototype branches; the resulting optimal values were relatively consistent across branches, and we selected rounded values that fell between the final tuned values of all three branches. These values were then held fixed in all subsequent experiments using contrastive regularization.
Prototype-based models underwent a three-stage training procedure (Section 3.5). Stage 1 (joint training) and Stage 2 (prototype projection) were repeated once to allow the prototypes to better converge on class-representative features, after which performance typically stabilized (no further increases in validation AUC). Stage 3 (fusion classifier training) was conducted only once, after freezing the prototype branches.
Appendix B System Requirements
All model training and inference were conducted on a high-performance computing cluster using a single NVIDIA A100 GPU with 40GB memory. Models were implemented in Python 3.10 using PyTorch 2.6.0 and PyTorch Lightning 2.5.0. Training visualization was performed with TensorBoard, and ECG data handling used Pandas and WFDB. Hyperparameter optimization was conducted with Optuna 4.2.1. All code was executed in a SLURM-managed Linux environment with CUDA 12 and cuDNN 9.1.
Appendix C ProtoECGNet Internal Architecture
Appendix D Example of a Case-Based Explanation from the 2D Global Branch
Figure 6 contains a case-based explanation for a diagnosis processed via the 2D global branch of ProtoECGNet.
Appendix E Prototype Review Interface
Figure 7: Clinician review interface for prototype evaluation.
Appendix F ECG and Prototype Visualization
Appendix G Diagnostic Groupings
Each of the 71 PTB-XL diagnostic labels was assigned to one of three prototype branches based on the type of visual reasoning required for diagnosis. These groupings were defined by two board-certified physicians (one cardiologist & one internist) and used to train the corresponding branch of ProtoECGNet. The full list of SCP codes and their assigned groupings is shown below.

### v5

@@ Appendix E Training Procedure
This stage used binary cross-entropy loss for multi-label prediction, with L1 regularization to encourage sparsity in the learned fusion weights. The sparsity constraint promoted more selective use of prototype information. Branch-specific classifiers were trained in the same way, to assess if the fusion classifier yielded performance benefits over simple macro-averaging. Additional training and tuning details are provided in Appendices F and G.
Appendix F Implementation Details
@@
Hyperparameter tuning was conducted using Optuna with 200 trials per model variant to optimize validation macro-AUROC. Tuned parameters included learning rate (range 1e-6 to 1e-2), weight decay (range 1e-6 to 1e-2), dropout rate (range 1e-6 to 1e-2), and learning rate scheduler (ReduceLROnPlateau, CosineAnnealingLR, or CyclicLR). Both the joint training phase and the classifier stage (including the fusion classifier and branch-specific classifiers) were tuned independently using this strategy. However, for the joint phase, we also tuned number of prototypes per class (range 1-20), and for the classifier stage, we tuned the L1 penalty (range 1e-6 to 1e-2). Our final model used five 1D rhythm prototypes, eighteen 2D partial prototypes, and seven 2D global prototypes per class.
When contrastive loss was not used, we adopted the prototype loss coefficients from Barnett et al. (2024). For contrastive models, we conducted a dedicated 1000-trial Optuna sweep to tune all four loss coefficients (range 1e-6 to 1e3). This sweep fixed non-loss hyperparameters (dropout = 0.3, batch size = 32, learning rate = 0.001, scheduler = ReduceLROnPlateau, L2 weight decay = $1\times 10^{-4}$), and only optimized the four loss weights. This was run separately for the 1D rhythm, 2D partial, and 2D global prototype branches; the resulting optimal values were relatively consistent across branches, and we selected rounded values that fell between the final tuned values of all three branches ($\lambda_{\text{clst}}=0.004,\lambda_{\text{sep}}=0.0004,\lambda_{\text{div}}=250,\lambda_{\text{cntrst}}=300$). These values were then held fixed in all subsequent experiments using contrastive regularization.
Appendix G System Requirements
All model training and inference were conducted on a high-performance computing cluster using a single NVIDIA A100 GPU with 40GB memory. Models were implemented in Python 3.10 using PyTorch 2.6.0 and PyTorch Lightning 2.5.0. Training visualization was performed with TensorBoard. Hyperparameter optimization was conducted with Optuna 4.2.1. All code was executed in a SLURM-managed Linux environment with CUDA 12 and cuDNN 9.1.
Appendix H Diagnostic Groupings
Each of the 71 PTB-XL labels was assigned to one of three prototype branches based on the type of visual reasoning required for diagnosis. These groupings were defined by two board-certified physicians (one cardiologist & one internist) and used to train the corresponding branch of ProtoECGNet. The full list of SCP codes and their assigned groupings is shown below.
@@ Appendix H Diagnostic Groupings
Appendix I Class-Specific Metrics
Class-specific AUROCs for the final contrastive ProtoECGNet fusion classifier are provided below. The number of samples of each class included in the PTB-XL test fold are included in parentheses after each class name. Interpret AUROCs for classes with less than 10 samples with caution. Confidence intervals were calculated by bootstrapping; they are “N/A” when only one value for a given class is present in the test fold.
NDT (182): 0.918 (0.900, 0.935)
NST (77): 0.851 (0.813, 0.887)
DIG (18): 0.936 (0.902, 0.966)
LNGQT (11): 0.834 (0.632, 0.981)
NORM (963): 0.930 (0.919, 0.940)
IMI (267): 0.890 (0.872, 0.907)
ASMI (234): 0.952 (0.939, 0.963)
LVH (214): 0.932 (0.915, 0.948)
LAFB (162): 0.967 (0.957, 0.976)
ISC (128): 0.956 (0.940, 0.969)
IRBBB (112): 0.924 (0.903, 0.943)
1AVB (79): 0.949 (0.933, 0.963)
IVCD (79): 0.656 (0.591, 0.718)
ISCAL (66): 0.937 (0.920, 0.952)
CRBBB (54): 0.997 (0.996, 0.999)
CLBBB (54): 0.996 (0.992, 0.999)
ILMI (48): 0.930 (0.880, 0.969)
LAO/LAE (42): 0.756 (0.680, 0.825)
AMI (35): 0.865 (0.810, 0.915)
ALMI (27): 0.957 (0.924, 0.981)
ISCIN (22): 0.928 (0.886, 0.963)
INJAS (22): 0.989 (0.981, 0.995)
LMI (20): 0.867 (0.804, 0.923)
ISCIL (18): 0.950 (0.921, 0.975)
LPFB (18): 0.935 (0.865, 0.984)
ISCAS (17): 0.972 (0.954, 0.985)
INJAL (14): 0.974 (0.942, 0.995)
ISCLA (13): 0.893 (0.832, 0.949)
RVH (12): 0.974 (0.957, 0.988)
ANEUR (10): 0.952 (N/A)
RAO/RAE (10): 0.951 (N/A)
EL (9): 0.836 (0.668, 0.951)
WPW (8): 0.881 (N/A)
ILBBB (8): 0.895 (N/A)
IPLMI (5): 0.772 (N/A)
ISCAN (4): 0.931 (N/A)
IPMI (3): 0.826 (N/A)
SEHYP (2): 0.989 (N/A)
INJIN (2): 0.996 (N/A)
INJLA (2): 0.800 (N/A)
PMI (2): 0.995 (N/A)
3AVB (2): 0.984 (N/A)
INJIL (2): 0.881 (N/A)
2AVB (1): 0.980 (N/A)
ABQRS (322): 0.781 (0.753, 0.808)
PVC (114): 0.980 (0.971, 0.987)
STD (101): 0.853 (0.818, 0.884)
VCLVH (87): 0.893 (0.864, 0.917)
QWAVE (55): 0.831 (0.781, 0.878)
LOWT (44): 0.904 (0.877, 0.929)
NT (42): 0.917 (0.886, 0.945)
PAC (40): 0.862 (0.803, 0.914)
LPR (34): 0.959 (0.943, 0.973)
INVT (29): 0.952 (0.929, 0.970)
LVOLT (18): 0.916 (0.871, 0.957)
HVOLT (6): 0.965 (N/A)
TAB (3): 0.864 (N/A)
STE (3): 0.669 (N/A)
PRC(S) (1): 0.867 (N/A)
SR (1674): 0.892 (0.875, 0.908)
AFIB (152): 0.982 (0.967, 0.993)
STACH (82): 0.981 (0.968, 0.990)
SARRH (77): 0.895 (0.862, 0.922)
SBRAD (64): 0.938 (0.901, 0.966)
PACE (28): 0.977 (0.942, 0.997)
SVARR (14): 0.829 (0.754, 0.900)
BIGU (8): 0.979 (N/A)
AFLT (7): 0.968 (N/A)
SVTAC (3): 0.995 (N/A)
PSVT (2): 0.999 (N/A)
TRIGU (2): 0.903 (N/A)